Lesson Thermodynamics Processes




1. An engine bought by an engineer rated a transfer of 5,000 J of energy from a hot reservoir during one cycle and dissipated 3x10³J. Find the efficiency of the engine.




2. A particular heat engine has a mechanical power output of 5.00 kW and an efficiency of 25.0%. The engine expels 8.00x10³ J of exhaust energy in each cycle. Find (a) the energy taken in during each cycle and (b) the time interval for each cycle

1
Expert's answer
2022-01-16T13:10:32-0500

1. The efficiency:


"\eta=\dfrac{5000-3000}{5000}·100\%=40\%"


2. (a) the energy taken in during each cycle:


"E_i=\dfrac{8000}{1-0.25}=10667\text{ J}."


(b) the time interval for each cycle:


"t=(E_i-E_e)/P=(10667-8000)/5000=0.53\text{ s}."